OpenStudy (anonymous):

I will give you a medal and be your fan What is the solution to the proportion? 3/m=12/15 A. m=2.4 B. m=3.75 C. m=4.5 D. m=4.75

OpenStudy (ash2326):

@jada1701877 Can you try cross multiplying? \[\frac 3 m =\frac{12}{15}\]

OpenStudy (anonymous):

Are you asking me if I know how to or can I do it with this problem?

OpenStudy (ash2326):

can you do that with this problem>

OpenStudy (anonymous):

I think so

OpenStudy (ash2326):

what will you get after cross multiplying?

OpenStudy (anonymous):

45

OpenStudy (anonymous):

I dont think you're suppose to croos multiply..in this problem

OpenStudy (ash2326):

\[\frac{3}{m}=\frac{12}{15}\] Cross multiply both sides \[45=12\times m\] Do you follow?

OpenStudy (anonymous):

Yess

OpenStudy (ash2326):

now divide both sides by 12 to find the value of m

OpenStudy (anonymous):

24

OpenStudy (ash2326):

\[12m=45\] \[12m\div 12= 45\div 12\] \[m=\frac{45}{12}\] Try again, 24 is not correct

OpenStudy (anonymous):

Well I did 45/12=3.75 @ash2326

OpenStudy (ash2326):

yes, you are correct!!

OpenStudy (anonymous):

Thanks so I just times then divide?

OpenStudy (ash2326):

yes, correct

OpenStudy (anonymous):

thank you